Feed aggregator

RF Interlock Faults not responding?

Recent Log Entries -

Log Number: 3856004Logbook(s): Entry Makers: kiarraBody: 

So, ideally, when rf drops from GDR to SEL mode the FSD should trip (i.e. stop beam production).
However, at this moment, FSD only trips when the RF gets turned OFF completely. Which
isn't very helpful for operators, or users, during production.

Auto-recovery would also be nice, but at this moment I personally don't mind the extra button clicking
to get beam back on target. My main concern is having the beam shut off on its own when machine status
is not what it should be -- especially since it's very easy to depolarize the HDIce target.

Is this an SSG thing? A Software thing?

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Recent Log Entries -

Log Number: 3856002Logbook(s): Body: 

(late entry) Day #29, Week 7 Day 4, Plans:

Tunnel --
* Gun2 Table Move (Mech Install, done)
* Gun3 stub, Y-chamber, NEG pedestal removals (Mech
Install, done)
* Wien cable pulls (Dela Cruz)

InjSB --
* PSS Upgrade, IN02B{NI, 29, 30} 9a - 4p (Beizer)
* Wien controls/HV/cable work, IN01B03, 6a - 1p (Dela
Cruz)
* Viewer work, IN03B07, 6a - 1p (Norris)
* BPM Work, IN03B{05-08, 23}, 5a - 1p (Francis)

Toolbox attended by -- Bullard, Hernandez-Garcia, Gould,
Croke, Grames, Cole, Mamun, Stutzman, Francis, Zhang,
Suleiman

Current UITF FSD Masking

Recent Log Entries -

Log Number: 3855998Logbook(s): Entry Makers: kiarraBody: 

A few HDIce components masked. And for some reason we're unable to unmask the
RF Quarter Cryomodule node.

Questions to address :

When will we be able to unmask the HDIce nodes? What exactly are they connected to/monitoring?

And what's going on with the RF node that's requiring the node to be masked?

I've attached the FSD node screens (fig 1) and the RF Combo screen / expert screens (fig 2-4).

todays UITF Day sshift plan

Recent Log Entries -

Log Number: 3855995Logbook(s): Entry Makers: poelkerBody: 

Today's DAY shift plan at UITF

1) harp swipe at the HDIce target using new harp
2) Kiarra and I will address FSD situations wityh goal of no masks iwhile operating,
3) setup alarm handler on wall monitor
4) after tune mode harp study, change fiber amp to support low current CW mode, giving beam to HDIce and letting them follow their strudy plan

iocea5 restored

Recent Log Entries -

Log Number: 3855990Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-10-15 09:37:04 The output from the myRestore execution was as follows: iocea5 Execute: /cs/opshome/scripts/60_second_wait.sh (success) Please wait 60 seconds for restore... Pause: 60 second(s) vac@2020-10-15 09:37 24(24) restored blmset@2020-10-15 09:37 24(24) restored system@2020-10-15 09:37 2(2) restored Execute: /cs/opshome/scripts/scan_gain_set.sh (success) Old : IPMBSCN Normal New : IPMBSCN Calib Old : IPMBSCN Cali

iocitf2 restored

Recent Log Entries -

Log Number: 3855988Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-10-29 09:28:33 The output from the myRestore execution was as follows: iocitf2 caen6533@2020-10-29 09:28 108(108) restored blm2@2020-10-29 09:28 48(48) restored systemEnum@2020-10-29 09:28 2(2) restored RasterUITF@2020-10-29 09:28 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ued -A -s -i iocitf2 (success) Procedure return status: Success

Pages

Subscribe to cebaf2013 aggregator